Let q(x)=5-x^2 , Evaluate and simplify the following:
q(3)=
q(r)=
q(r-3)=
q(r)-3=
q(r)-q(3)=

Answer by Cromlix(4381)

Hi there,
Let q(x)=5-x^2 , Evaluate and simplify the following:
q(3)= 5 - (3)^2 = 5 - 9 = -4
q(r)= 5 -(r)^2 = 5 - r^2
q(r-3)= 5 - (r - 3)^2 = 5 - (r^2 - 6r + 9) = 5 - r^2 + 6r - 9
= - r^2 + 6r - 4
q(r)-3= 5 - r^2 - 3 = 2 - r^2
q(r)-q(3)= (5 - r^2)- (-4) = 5 - r^2 + 4 = 9 - r^2
Hope this helps :-)